Expectations of Functions:

Let X be distributed over the set N of non-negative integers, with probability mass function:

Let X be distributed over the set N of non-negative integers, with pmf a P(X = i) = 21 for some fixed α E R. Find EX For Y X mod 3, find ·P(Y= 1) .ELY
Answer #1

for above to be valid:

1)\sumP(x) =1

here \sumP(x) =\alpha*(1/20+1/21+1/22+1/23+...) =\alpha*(1/(1-1/2)) =1

\alpha =1/2

2) E(X)=\sumxP(x) =(1/2)*(0*1/20+1/21+2/22+3/23+...) =(1/2)*(1/2)/(1-1/2)2 =1

3)P(Y=1) =P(X=1)+P(X=4)+P(X=7)+P(X=8)+..... =(1/2)2+(1/2)5+(1/2)8+.... =(1/4)/(1-1/8)=2/7

4)

E(Y)=0*(P(X=0)+P(X=3)+P(X=6)+...)+1*(P(X=1)+P(X=4)+P(X=7)+..)+2*(P(X=2)+P(X=5)+P(X=8)+..)

=1*(2/7)+2*(1/7)=4/7